A child runs 5 m/s down an airport walkway that is moving at 2 m/s in the same direction. What distance does the child cover in 10 seconds?
Answer #1

Effective speed = 5+2 = 7 m/s

Time t = 10 s

We know,

D = vt

= 7 * 10

= 70 m (answer)